# Collaborative Unlabeled Data Optimization

### v1

@@ 4 Experiments
In this section, we outline the experimental setup in Section 4.1 and assess the performance of our proposed collaborative data optimization framework, CoOpt, across various real-world scenarios, employing different datasets and architectures, as discussed in Section 4.2. Subsequently, we explore a continuous data optimization framework that allows the prior models of participants to evolve, thereby further optimizing the targets, as detailed in Section 4.3. Finally, we validate the effectiveness of the introduced uniform value metric and the target alignment strategy in Section 4.4.
@@ 4.1 Experimental Setup
We conduct experiments on both large-scale and small-scale datasets, including the full $224\times 224$ ImageNet-1k (224 $\times$ 224) [10], Tiny-ImageNet (64 $\times$ 64) [21], CIFAR-100 [19] and CIFAR-10 (32 $\times$ 32) [20]. Following previous self-supervised studies [15, 6, 13, 8, 1, 40], we employ a range of model capacities backbone architectures to evaluate the generalizability of our method, including ResNet-18, 50, 101 [16], ViT [12], and a series of CLIP-based models [26].
@@
For the unlabeled data, referring to a prior widely-used benchmark [9], we consider several state-of-the-art self-supervised methods as baselines for a broader practical impact, including: SimCLR [6], BYOL [13], DINO [4], MoCo [15], SimSiam [8], SwAV [3], and DCL [36]. Notably, we do not compare our approach with knowledge distillation (KD) or dataset distillation (DD) methods, as summarized in Table 1, since they primarily foucus on the optimization of labeled data. In contrast, this work focuses on the utilization of unlabeled data.
@@
Following previous benchmarks [15, 6, 13, 8], we evaluate the representation quality of the trained models by evaluating their test accuracy (%) using an offline linear probing strategy. Additionally, computational efficiency quantified by time cost (s).
@@
The proposed algorithm, CoOpt, involves an open data platform and facilitates interaction among multiple participants. The implementation follows five key steps (detailed in Section 3.2), and more details are provided in Appendix B.
@@
    For steps 2 and 3, in practical applications, each participant can use publicly pre-trained models or their own models directly as the prior model. To simulate the diversity of prior models in practical applications, we utilize pre-trained models trained on ImageNet as well as a series of pre-trained CLIP-based models.
@@
4.2 Comparison with the state-of-the-art methods
We evaluate our framework, CoOpt, across various scenarios: (1) participants leverage a diverse array of pre-trained prior models, characterized by variations in datasets, architectures, and training paradigms. (2) we specifically evaluate the robustness of our method when employing different prior datasets used for training prior models; (3) finally, we explore special cases involving human or weak models, particularly where there are resource-rich or resource-poor participants.
Diverse Pre-trained Prior Models.
Given the diversity of prior models, we employ a comprehensive set of pre-trained models, including those trained on ImageNet-1K and CLIP-based models. Details of these models are provided in Appendix B.2. These models are trained using various training paradigms, datasets, and architectures. As shown in Table 2, our method CoOpt demonstrates superior performance and training efficiency compared to existing self-supervised learning methods. Additional experimental results for the optimized data, trained across a wider range of architectures, are presented in Table 5 within Appendix B.3.
  1. (a)
    Performance: CoOpt significantly outperforms state-of-the-art self-supervised learning methods. For instance, it achieves an improvement of 13.6% and 6.8% over the leading self-supervised approach BYOL on Tiny-ImageNet and ImageNet-1K, respectively. Notably, CoOpt demonstrates more significant improvements on larger-scale datasets, such as Tiny-ImageNet and ImageNet-1K, which is particularly advantageous given the current emphasis on large-scale data era.
  2. (b)
@@ 4.2 Comparison with the state-of-the-art methods
    Efficiency: CoOpt demonstrates a substantial reduction in training costs across various datasets. Notably, on the Tiny-ImageNet, CoOpt achieves a training speed that surpasses BYOL by a factor of approximately ×$\times 1.94$.
@@
Dataset BYOL (Baseline) Our CoOpt (Diverse Prior Datasets)
@@
Diverse Prior Datasets.
In the last experiment, we employ pre-trained models, some of which are congruent with the unlabeled training data, while others are incongruent. We refer to these datasets as the prior datasets. To rigorously evaluate the influence of these prior datasets, we perform an analysis across scenarios where the prior datasets used for prior models either align with or differ from the unlabeled training dataset. For instance, in the aligned scenario, the training dataset is CIFAR-10, and the prior models are also trained on CIFAR-10 (CIFAR-10 (P)). Conversely, in the divergent scenario, the training dataset remains CIFAR-10, while the prior models are trained on CIFAR-100 (P). Details of these models are provided in Appendix B.4. We evaluate our approach on four publicly available datasets, with the results summarized in Table 3.
  1. (a)
    Utilizing prior models trained on the same dataset as the target training data can lead to notable performance gains.
  2. (b)
    For complex training datasets, leveraging prior models trained on simpler datasets may degrade performance compared to BYOL This degradation arises due to the limited informativeness of simpler prior datasets, which provide weaker guidance.
  3. (c)
    Importantly, for all unlabeled training datasets, employing prior models trained on ImageNet-1K consistently yields significant performance gains. This can be attributed to the strong generalization abilities of such models. This observation is particularly pertinent in practical applications, as the majority of publicly available pre-trained models are derived from ImageNet-1K.
@@
CoOpt 88.7 $\pm$ 0.1
89.0 $\pm$ 0.0
90.3 $\pm$ 0.0
89.2 $\pm$ 0.0
Extreme Cases of Prior Models: Human or Weak.
@@
In real-world applications, extreme cases arise due to the varying capabilities of participants. For example, some participants have extensive resources and employ human annotators for labeling, while others may have limited resources and rely on weak models with inferior performance and generalization abilities. In our experiments, we define weak models as those trained during intermediate stages that are even far from convergence. To simulate the conditions, in addition to the prior models used in the first experiment, we incorporate $5$ prior models, either human or weak models to the data optimization process. The results are summarized in Table 4. Surprisingly, even the inclusion of weaker models contributes to enhancing the final performance, indicating that such models can still provide valuable information. Moreover, it is important to note that the integration of high-capacity, human-like models results in significant performance improvements.
(a) Training curves
(b) Varying Prior Models
(c) Uniform value
(d) Shared data size
Figure 3: Comprehensive Analysis of CoOpt. (a) Training curves: Comparison of various self-supervised learning methods and the proposed CoOpt. (b) Prior Models With Varying Accuracies. Employing a strong model as the prior model is not essential (c) Correlation Verification: Examine the relationship between the uniform value and performance, which demonstrates a strong negative correlation. (d) Influence of shared data size: As the size of the shared data increases, the performance gains diminish.
Moreover, we conduct experiments with only weaker model, as shown in Figure 3(b). Here, The dashed line represents the training curve of BYOL (baseline), while the solid lines correspond to prior models with different accuracies. "Prior model (BYOL)" indicates the use of BYOL as the prior model, and the accuracies of the other prior models are all lower than that of BYOL. Notably, most prior models surpass BYOL performance in the early stages of training, and even a model with approximately 75% accuracy can outperform BYOL in the final performance. The results demonstrate that employing a strong model as the prior model is not essential.
Figure 4: A Practical Scenario: Continuous Optimization Prior models experience temporal evolution, leading to improved target quality across multiple rounds.
4.3 Continuous Data Optimization
We explore another practical scenario where the prior models undergo temporal evolution. For example, a participant’s initial model, such as ResNet-50, might be upgraded to a higher-capacity model like ResNet-101 as their computational resources improve. Consequently, the process can be treated as a dynamic and continuous procedure. The detailed description of the process is provided in Appendix B.5.
We simulate this scenario by conducting experiments where, in each interaction round between the platform and participants, $20$% of the participants randomly update their model architectures to reflect an increase in model capacity. The training curves across 10 rounds on CIFAR-100 are depicted in Figure 4. The results have demonstrated that in CoOpt, as the prior models evolve, the quality of the targets improves, thereby facilitating continuous optimization. In particular, compared to the baseline BYOL, these improvements can achieve an enhancement of $4.6$%.
4.4 Ablation Study
Effectiveness of Uniform Value.
To evaluate the effectiveness of uniform value in estimating the quality of prior models, we employ a diverse set of prior models, calculating both their uniform value and test accuracy. We employ the Spearman rank correlation coefficient [^4] The formula is: $\rho=1-\frac{{6\sum d_{i}^{2}}}{{n(n^{2}-1)}}$, where $d_{i}$ represents the difference between the ranks of each pair of observations and $n$ denotes the number of observations. $\rho$ [39] to quantify the association between Uniform Value and accuracy. As illustrated in Figure 3(c), the Spearman rank correlation coefficient is $\rho=-0.9714$, indicating a strong correlation between uniform value and model performance, thereby effectively assessing the quality of prior models.
Target Distribution Alignment.
In the real-world applications, participants often employ diverse prior models, which results in target distribution space inconsistencies, as demonstrated in Figure 5(a). We conduct an ablation study w/o alignment and w/ alignment to demonstrate the importance and necessarily of target distribution space alignment, and the training curves presented in Figure 5(d). Three key observations include:
  1. (a)
    Compared to without alignment (green line), our proposed approach (blue line) yields a performance improvement of $16.9$%, which highlights the critical importance of aligning the target distribution space. Notably, the training curve without alignment (green line) initially ascends before declining, suggesting that during the early phase of model training, optimizing the target leads to performance enhancement. However, as training continues, the severely inconsistent targets significantly degrade model performance.
  2. (b)
    Furthermore, we compare our alignment strategy, which aligns to the best prior model, with two other straightforward strategies: aligning to a medium prior model (purple line) and aligning to the worse prior model (red line). Obviously, three strategies outperform the scenario without alignment, and aligning to the best prior model provides the most significant performance gains.
  3. (c)
    To further analyze the underlying reasons, we employ t-SNE visualization. Notably, a comparison between alignment with the better participant ( Figure 5(b)) and the worse participant ( Figure 5(c)) reveals that alignment with a worse-quality model diminishes the representative capability of the targets. This, in turn, results in less effective guidance for model training.
Influence of Shared Unlabeled Data Size.
The shared unlabeled data $S$ is used to estimate the uniform value and compute the transformation matrix for target alignment, as detailed in Section 3.4. To explore the influence of the size of shared unlabeled data, we conduct experiments on both CIFAR-10 and CIFAR-100, varying the size of the shared data in proportions of $\{0.01,0.05,0.1,0.2,0.4,0.6,0.8\}$ relative to the original dataset. The results are presented in Figure 3(d). It is observed that as the size of the shared data increases, performance gains become marginal. Specifically, increasing the proportion of data from $0.01$% to $0.05$% results in improvements of $2.4$% and $2.5$% for CIFAR-10 and CIFAR-100, respectively. However, increasing the proportion from $0.1\%$ to $0.8\%$ results in only a $0.5\%$ and $0.6\%$ improvement. This suggests that a small-sized shared unlabeled dataset is adequate for accurate uniform value estimation and the computation of the transformation matrix.
@@ 4.4 Ablation Study
5 Conclusion and Future Work
Conclusion.

### v2

@@ 4 Experiments
We conduct extensive experiments to demonstrate the key advantages of CoOpt in Sec. 4.2. Specifically, first, to evaluate its efficacy and efficiency in utilizing unlabeled data, we compare CoOpt with state-of-the-art self-supervised learning methods. Second, to examine the necessity of distributed optimization, we further compare CoOpt with centralized optimization approaches. Third, we train diverse model architectures on the optimized data, thereby assessing its generalizability and reusability. Forth, we demonstrate the potential of CoOpt for continuous data optimization, showing how it continuously enhances data quality. Furthermore, we explore factors that influence its effectiveness (Sec. 4.3), including different prior datasets and prior models. Finally, we present comprehensive ablation studies (Sec. 4.4) to verify the impact of each module in CoOpt and its lightweight design.
@@ 4.1 Experimental Setup
We conduct experiments on both large-scale and small-scale datasets, including ImageNet-1k (224 $\times$ 224) (Deng et al., 2009), Tiny-ImageNet (64 $\times$ 64) (Le & Yang, 2015), CIFAR-100 (Krizhevsky et al., 2009a) and CIFAR-10 (32 $\times$ 32) (Krizhevsky et al., 2009b). Following previous self-supervised studies (He et al., 2020; Chen et al., 2020a; Grill et al., 2020; Chen & He, 2021; Assran et al., 2023; Zhang et al., 2024), we employ a range of backbone architectures to evaluate the generalizability of our method, including ResNet-{18, 50, 101} (He et al., 2016), ViT (Dosovitskiy et al., 2020), and a series of CLIP-based models (Radford et al., 2021).
@@
For the unlabeled data, following a widely used benchmark (Da Costa et al., 2022), we compare against state-of-the-art self-supervised (SSL) methods, including: SimCLR (Chen et al., 2020a), BYOL (Grill et al., 2020), DINO (Caron et al., 2021), MoCo (He et al., 2020), SimSiam (Chen & He, 2021), SwAV (Caron et al., 2020), and DCL (Yeh et al., 2022). Notably, we do not compare with knowledge distillation (KD) or dataset distillation (DD) methods, since the training paradigm of KD differs significantly from ours, while DD primarily focuses on the labeled data.
@@
Following previous benchmarks (Grill et al., 2020; Chen & He, 2021), we evaluate the representation quality of models by evaluating their test accuracy (%) using an offline linear probing strategy. Additionally, computational efficiency quantified by time cost (s).
@@
The proposed algorithm, CoOpt, involves an open data platform and facilitates interaction among multiple participants. The implementation follows five key steps (detailed in Sec. 3.2), and more details are provided in App. B.
@@
    For steps 2 and 3, in practical applications, each participant can use publicly pre-trained models or their own models directly as the prior model. To simulate the diversity of prior models in practical applications, we use a series of pre-trained CLIP-based models.
@@
4.2 What are the advantages of CoOpt?
Dataset Metric BYOL DINO MoCo SimCLR SimSiam SwAV DCL CoOpt (Ours)
CF-10 Acc. (%) 82.8 $\pm$ 0.1 82.6 $\pm$ 0.0 82.9 $\pm$ 0.1 83.1 $\pm$ 0.0 79.0 $\pm$ 0.0 82.9 $\pm$ 0.1 83.9 $\pm$ 0.1 89.5 $\pm$ 0.1 ($\uparrow$ 5.6)
CF-100 Acc. (%) 51.7 $\pm$ 0.1 51.0 $\pm$ 0.0 57.8 $\pm$ 0.1 55.4 $\pm$ 0.0 44.6 $\pm$ 0.1 53.2 $\pm$ 0.1 58.2 $\pm$ 0.2 67.3 $\pm$ 0.1 ($\uparrow$ 9.1)
T-IN Acc. (%) 43.9 $\pm$ 0.2 36.1 $\pm$ 0.0 42.4 $\pm$ 0.2 41.5 $\pm$ 0.1 40.8 $\pm$ 0.0 39.9 $\pm$ 0.1 44.6 $\pm$ 0.0 60.3 $\pm$ 0.1 ($\uparrow$ 15.7)
IN-1K Acc.(%) 61.9 $\pm$ 0.1 52.2 $\pm$ 0.0 57.6 $\pm$ 0.0 58.0 $\pm$ 0.0 55.8 $\pm$ 0.1 57.2 $\pm$ 0.1 60.6 $\pm$ 0.1 69.8 $\pm$ 0.1 ($\uparrow$ 7.9)
Comparison with SSL Methods.
@@ 4.2 What are the advantages of CoOpt?
As shown in Tab. 1, our CoOpt demonstrates superior performance and efficiency compared to existing self-supervised learning methods. We also visualize the training dynamic in Fig. 5(a). Specifically, CoOpt achieves an improvement of 7.9% over the leading self-supervised approach BYOL on ImageNet-1K. For Efficiency, CoOpt demonstrates a substantial improvememt across various datasets. Notably, on the Tiny-ImageNet, CoOpt achieves a training speed that surpasses the efficient method SimSiam by a factor of approximately $\times 1.94$.
Table 2: Comparison with Centralized Optimization.
Method Time (s) Acc. (%)
Centralized 23.71 62.1 $\pm$ 0.1
Ours 16.31 65.8 $\pm$ 0.1
Comparison with Centralized Optimization.
A key advantage of CoOpt lies in its ability to use diverse prior models in parallel, thereby enabling efficient optimization. To validate this, we compare CoOpt with centralized optimization, where a single model is used to optimize all unlabeled data. Specifically, we consider 10 prior models of varying quality on CIFAR-100. For the centralized setting, we report the mean performance of these 10 independently selected models to ensure fairness. As summarized in Tab. 2, CoOpt consistently demonstrates superior efficiency and efficacy, verifying the benefits of distributed over centralized optimization. This improvement arises from the proposed target alignment strategy (Sec. 3.4), which leverages high-quality priors to enhance the target distribution of weaker models.
While one might envision an ideal centralized solution that exclusively employs the best prior model, such an approach is rarely practical in real-world scenarios A major concern is fairness and computational burden, as concentrating all computation on a single party imposes excessive cost and discourages participation. Another challenge is privacy, since high-performing models are typically proprietary, and centralizing them may violate ownership or data-sharing constraints. In contrast, CoOpt collaboratively leverage diverse prior models, achieving competitive performance while substantially reducing the burden on any individual participant.
Table 3: Comparison of CoOpt with BYOL on Diverse Networks.
Method
Network BYOL CoOpt
ResNet-50 60.4 $\pm$ 0.1 63.8 $\pm$ 0.0
ResNet-101 61.5 $\pm$ 0.2 65.7 $\pm$ 0.2
MobileNet-v2 24.0 $\pm$ 0.5 58.1 $\pm$ 0.0
Efficientnet-b0 2.3 $\pm$ 0.2 70.7 $\pm$ 0.2
ViT 38.5 $\pm$ 0.1 57.8 $\pm$ 0.1
Generalizability and Reusability of Optimized Data.
Another key advantage of our optimized data lies in its strong generalizability and reusability: once constructed, it can be directly employed for downstream diverse model training without further modification. To evaluate this advantage, we conduct experiments by training a variety of neural architectures on the optimized data and compare the results to a strong baseline, BYOL, which relies on training from scratch on the original unlabeled dataset. The results are summarized in Tab. 3.
Obviously, CoOpt consistently delivers significant performance improvements over BYOL across multiple architectures. In particular, BYOL suffers substantial degradation when applied to lightweight networks such as MobileNet-v2. A plausible explanation is that these models are more sensitive to unstable batch normalization (BN) statistics in early network layers. Instead, our optimized data exhibits strong generalization across diverse architectures.
Figure 4: A Practical Scenario: Continuous Optimization.
Continuous Data Optimization.
We further explore a practical scenario where the prior models undergo temporal evolution. For example, a participant’s initial model, such as ResNet-50, might be upgraded to a higher-capacity model like ResNet-101 as their computational resources improve. Consequently, the process can be treated as a dynamic and continuous procedure. The detailed description of the process is provided in App. B.4.
We simulate this scenario by making random 2020% of the participants increase their model capacity in each round. The training curves across 10 rounds on CIFAR-100 are shown in Fig. 4. The results demonstrate that in CoOpt, as the prior models evolve, the quality of the targets improves, thereby facilitating continuous optimization. In particular, over 10 rounds, the continuous optimization setting yields a 4.6% performance gain.
4.3 What influence the effectiveness of CoOpt?
We investigate the key factors that influence the effectiveness of the optimized data. Since the optimization of unlabeled data in CoOpt relies solely on task-agnostic prior models, we focus on two primary aspects in these experiments: prior datasets and prior models.
(a) Training Curves
(b) Varying Prior Models
(c) Uniform Value
(d) Shared Data Size
Figure 5: Comprehensive Analysis of CoOpt. (a) Training curves: Comparison of SSL methods and our CoOpt. (b) Prior Models With Varying Accuracies. Even with a very weak prior model, CoOpt accelerates the early-stage training. (c) Correlation Verification: Verify the correlation between the uniform value and performance. (d) Influence of shared data size: As shared data’s size increases, the performance gains diminish.
@@ 4.3 What influence the effectiveness of CoOpt?
Dataset
BYOL
Our CoOpt (Diverse Prior Datasets)
@@
Influence of Prior Datasets.
To rigorously evaluate the influence of prior datasets that are used for training prior models, we perform an analysis across scenarios where the prior datasets used for prior models either align with or differ from the unlabeled training dataset. For instance, in the aligned scenario, the training dataset is CIFAR-10, and the prior models are also trained on CIFAR-10 (CIFAR-10 (P)). Conversely, in the divergent scenario, the training dataset remains CIFAR-10, while the prior models are trained on CIFAR-100 (P). Details of these models are provided in App. B.3. We evaluate our approach on four publicly available datasets, with the results summarized in Tab. 4.
Obviously, for all unlabeled training datasets, employing prior models trained on ImageNet-1K consistently yields notable performance gains, attributed to their strong generalization abilities. This observation is particularly relevant in practical applications, as most publicly available pre-trained models are derived from ImageNet-1K or even larger datasets. On the other hand, for complex training datasets, leveraging prior models trained on simpler datasets may result in degraded performance compared to BYOL. This is likely due to the limited informativeness of simpler prior datasets, which provide weaker guidance. We further examine the influence of weak models in Fig. 5(b).
@@
CoOpt 89.5 $\pm$ 0.1
89.2 $\pm$ 0.2
90.5 $\pm$ 0.1
89.8 $\pm$ 0.1
Special Cases of Prior Models: Human or Weak Involvement.
@@
In real-world applications, extreme cases arise due to the varying capabilities of participants. For example, some participants have extensive resources and can employ human annotators for labeling, while others may have limited resources and rely on weak models with inferior generalization abilities. In this experiment, we define weak models as those trained during intermediate stages that are even far from convergence. To simulate the conditions, in addition to the prior models used in the first experiment, we incorporate $5$ prior models, either human or weak models to the data optimization process. The results are summarized in Tab. 5. Surprisingly, even the inclusion of weaker models contributes to enhancing the final performance, indicating that such models can still provide valuable information. Moreover, it is important to note that the integration of high-capacity, human-like models results in significant performance improvements.
@@
Extreme Cases of Prior Models: All are Weak.
Moreover, we conduct experiments with only weak model, as shown in Fig. 5(b). Here, the dashed line represents the training curve of BYOL (baseline), and the solid lines means prior models with different accuracies. "Prior model (BYOL)" indicates the use of BYOL as the prior model, and the accuracies of the other prior models are all lower than that of BYOL. While a stronger prior model does yield better performance, our results demonstrate that even with a moderately weak prior model with approximately 75% accuracy, our method can still outperform the baseline. More importantly, even when prior models are substantially weak, CoOpt still significantly outperform baseline in the early training stages.
4.4 Ablation Study
Can Uniform Value Effectively Assess the Quality of Prior Models?
To evaluate the effectiveness of uniform value in estimating the quality of prior models, we employ a diverse set of prior models and compute both their uniform values and corresponding test accuracies. To quantify the relationship between these two measures, we adopt the Spearman rank correlation coefficient [^3] The formula is: $\rho=1-\frac{{6\sum d_{i}^{2}}}{{n(n^{2}-1)}}$, where $d_{i}$ represents the difference between the ranks of each pair of observations and $n$ denotes the number of observations. $\rho$ (Zar, 2014) to quantify the association between Uniform Value and accuracy. As illustrated in Fig. 5(c), the Spearman rank correlation coefficient is $\rho=-0.9714$, indicating a strong correlation between uniform value and model performance, thereby effectively assessing the quality of prior models.
Why Target Distribution Alignment is Necessary?
In real-world applications, participants often adopt diverse prior models, leading to inconsistencies in the target space (Fig. 6(a)). To verify the necessity of alignment, we conduct an ablation study with and without alignment, and the training curves in Fig. 6(d) show that our alignment method improves performance by 16.9%, veifying its effectiveness. Furthermore, We compare three strategies: aligning to the best prior model (ours), a medium-quality prior model (purple line), and the worst prior model (red line). All strategies outperform the unaligned baseline, but aligning to the best prior model yields the largest performance gains, as reported in Tab. 7 of App. B.5. To analyze the underlying reasons, we employ t-SNE visualization. As shown in Fig. 6(b) and Fig. 6(c), alignment with a high-quality model enhances the representational capability of the targets, whereas alignment with a poor model diminishes it.
How Much Shared Unlabeled Data Is Enough?
The shared unlabeled data $S$ is used to estimate the uniform value and compute the transformation matrix for target alignment, as detailed in Sec. 3.4. To explore the influence of the data size, we vary the proportion of shared data on CIFAR-10 and CIFAR-100, as shown in Fig. 5(d). Obviously, as the size of the shared data increases, performance gains become marginal, indicating that a very small fraction (around 0.05%) is sufficient for accurate estimation and alignment. We further validate this on ImageNet-1K (Tab. 8 in App. B.6), where only 0.001% of the data achieves comparable results to larger proportions. This minimal requirement imposes negligible additional overhead, ensuring scalability to very large datasets.
Table 6: Comparison of Time (s) on ImageNet-1K.
BYOL Uniform value Alignment
133,766.19 139.16 36.97
Does Alignment Incur Significant Overhead?
We report the computational cost of uniform value estimation and alignment in Tab. 6, verifying that both incur only negligible cost. This efficiency stems from the former requiring just a single forward pass to obtain targets, while the latter involves optimizing a lightweight matrix.
5 Conclusion